Required information

[The following information applies to the questions displayed below.]  

Mom’s Apple Pie Company uses a standard cost system. The standard direct labor time for each pie is 10 minutes. During the most recent month, the company produced and sold 6,000 pies. The standard direct labor rate is $8 per hour; the actual labor rate per hour for the month was $8.40. The company used a total of 980 labor hours.

What was the direct labor rate variance for the month? Was this variance favorable (F) or unfavorable (U)? (Do not round intermediate calculations.)
Answer #1

Labor rate variance

= (SR-AR) * AH

= (8 - 8.40) * 980

= 0.40 * 980

= 392 U

Comment if you face any issues